OpenStudy (anonymous):

Jar B has 20% more marbles than jar A. What percent of the marbles in jar B have to be moved to jar A, in order for the number of marbles in both jars to be equal?

OpenStudy (anonymous):

try your sense

OpenStudy (anonymous):

what?

OpenStudy (phi):

Try have A contain 100 (nice number!) B has 20% more: 20% is 20/100 or 0.20 20% of 100 = 0.02*100= 20 so B has 100+20 = 120 to make the number equal, it looks like you want to move 10 from B to A so both have 110 what percent is 10/120 ?

OpenStudy (anonymous):

for every 100a you'll have 120 b. so you need to move 10 toa for every 100, so 10%

OpenStudy (phi):

@sadrian I would seem not. Try will some example numbers. for example 120 in B, 100 in A

OpenStudy (anonymous):

@phi yup i misread.